# Encyclo-K: Evaluating LLMs with Dynamically Composed Knowledge Statements

### v1

@@ Abstract
Benchmarks play a crucial role in tracking the rapid advancement of large language models (LLMs) and identifying their capability boundaries. However, existing benchmarks predominantly curate questions at the question level, suffering from three fundamental limitations: vulnerability to data contamination, restriction to single-knowledge-point assessment, and reliance on costly domain expert annotation. We propose Encyclo-K, a statement-based benchmark that rethinks benchmark construction from the ground up. Our key insight is that knowledge statements, not questions, can serve as the unit of curation, and questions can then be constructed from them. We extract standalone knowledge statements from authoritative textbooks and dynamically compose them into evaluation questions through random sampling at test time. This design directly addresses all three limitations: the combinatorial space is too vast to memorize, and model rankings remain stable across dynamically generated question sets, enabling reliable periodic dataset refresh; each question aggregates 8–10 statements for comprehensive multi-knowledge assessment; annotators only verify formatting compliance without requiring domain expertise, substantially reducing annotation costs. Experiments on over 50 LLMs demonstrate that Encyclo-K poses substantial challenges with strong discriminative power. Even the top-performing OpenAI-GPT-5.1 achieves only 62.07% accuracy, and model performance displays a clear gradient distribution—reasoning models span from 16.04% to 62.07%, while chat models range from 9.71% to 50.40%. These results validate the challenges introduced by dynamic evaluation and multi-statement comprehensive understanding. These findings establish Encyclo-K as a scalable framework for dynamic evaluation of LLMs’ comprehensive understanding over multiple fine-grained disciplinary knowledge statements.
@@ 1 Introduction
As large language models (LLMs) advance rapidly, benchmarks for language understanding evolve accordingly. Early benchmarks such as GLUE [DBLP:conf/iclr/WangSMHLB19] and SuperGLUE [DBLP:conf/nips/WangPNSMHLB19] play a pivotal role in advancing fundamental language understanding tasks. To keep pace with improving model capabilities, subsequent benchmarks including MMLU [hendrycks2020measuring], MMLU-Pro [wang2024mmlupro], TriviaQA [joshi2017triviaqa0], and HotpotQA [yang2018hotpotqa0] raise the bar by broadening knowledge coverage and increasing task complexity. More recently, GPQA [rein2023gpqa], SuperGPQA [pteam2025supergpqascalingllmevaluation], and HLE [phan2025humanity0s] push the frontier further by demanding expert-level knowledge comprehension and deep reasoning.
Despite this progress, these benchmarks predominantly collect questions that already exist in textbooks or on websites. This paradigm of curating benchmarks at the question level suffers from three key limitations. First, questions or their variants may already appear in model training corpora. Even if current data contamination is absent, future training runs remain at risk, potentially distorting evaluation results. Second, each question typically targets a single knowledge point, making it difficult to assess comprehensive understanding across multiple knowledge points. Third, curating high-quality questions requires domain experts to design, annotate, and review them, incurring substantial cost and time.
@@
Figure 1: Pipeline of Encyclo-K.
Concretely, as illustrated in Figure 1, we extract standalone knowledge statements from authoritative textbooks and compose them into evaluation questions through random sampling at test time. This simple pipeline directly addresses all three limitations. For contamination resistance, we randomly sample and combine statements into questions at each test time. Even if individual statements appear in training data, their compositions form a combinatorial space too vast to memorize. For comprehensive assessment, each question aggregates 8–10 statements and asks models to identify which are correct. This requires joint understanding across diverse statements, going beyond what single-statement questions can probe. For cost reduction, correct statements are extracted directly from textbooks while incorrect ones are generated by reasoning models. Annotators only verify that each statement is self-contained and clearly written, requiring no domain expertise.
Extensive experiments on over 50 LLMs validate the effectiveness of Encyclo-K. Regarding contamination resistance, model performance remains highly stable across five dynamically generated question sets with different random seeds, demonstrating that Encyclo-K can reliably release multiple evaluation sets to prevent overfitting. Joint comprehension of multiple statements poses substantial challenges: even state-of-the-art models like OpenAI-GPT-5.1 achieve only 62.07% accuracy, with performance degrading systematically as option count increases from 4 to 10.
@@
  • We rethink benchmark construction by treating individual knowledge statements, rather than questions, as the atomic unit of curation. This shift enables diverse sourcing from authoritative materials, eliminates the need for domain-expert annotation, and substantially reduces construction costs.
  • We compose statements into questions through random sampling, yielding two key properties: (1) Dynamic—the combinatorial space is too vast to memorize, with experiments across multiple random seeds confirming stable model rankings; (2) Multi-statement—each question requires joint comprehension of multiple statements, posing greater challenges than single-knowledge-point questions.
@@ 2.2 LLMs Evaluation Benchmarks
To ensure a robust evaluation of model generalizability and alleviate question leakage, we designed two distinct mechanisms. First, Encyclo-K supports dynamic evaluation: statements within questions can be substituted, making the question content itself dynamically variable and enabling continuous generation of novel test samples to effectively address data contamination. Second, each question integrates multiple knowledge, emphasizing comprehensive understanding while ensuring that individual questions better represent the overall knowledge structure of a discipline.
3 DynamicGPQA Benchmark Dataset
@@ 5.1 Main Results
We first compare reasoning models against their chat counterparts within the same model family. As shown in Table 3, Qwen3-30B-A3B-Thinking (40.57%) outperforms Qwen3-30B-A3B-Instruct (36.72%) by 4 percentage points, and Qwen3-4B-Thinking (33.45%) exceeds Qwen3-4B-Instruct (22.55%) by 11 percentage points. While this comparison suggests reasoning benefits, it conflates reasoning capability with potential differences in training data or optimization objectives between model variants.

### v2

@@ Abstract
Robust benchmarks are essential for accurately reflecting the generalization capabilities of large language models (LLMs). Existing benchmarks that curate questions at the question level suffer from three limitations: vulnerability to data contamination, restriction to single-concept assessment, and reliance on costly domain expert annotation. We propose Encyclo-K, a statement-based benchmark that extracts standalone knowledge statements from authoritative textbooks and dynamically composes them into evaluation questions through random sampling at test time. This design directly addresses all three limitations: the combinatorial space resists memorization while maintaining stable model rankings across question sets; each question aggregates 8–10 statements for comprehensive knowledge assessment; and annotators only verify formatting compliance without requiring domain expertise. Experiments on over 50 LLMs demonstrate that Encyclo-K poses substantial challenges—even OpenAI-GPT-5.1 achieves only 62.07% accuracy, with model performance displaying clear gradient distributions across both reasoning models (16.04%–62.07%) and chat models (9.71%–50.40%). Encyclo-K achieves robust LLM evaluation through contamination-resistant dynamic generation and comprehensive multi-statement assessment.
@@ 1 Introduction
As LLMs advance rapidly, benchmarks for language understanding evolve accordingly. Early benchmarks such as GLUE [DBLP:conf/iclr/WangSMHLB19] and SuperGLUE [DBLP:conf/nips/WangPNSMHLB19] play a pivotal role in advancing fundamental language understanding tasks. To keep pace with improving model capabilities, subsequent benchmarks including MMLU [hendrycks2020measuring], MMLU-Pro [wang2024mmlupro], TriviaQA [joshi2017triviaqa0], and HotpotQA [yang2018hotpotqa0] raise the bar by broadening knowledge coverage and increasing task complexity. More recently, GPQA [rein2023gpqa], SuperGPQA [pteam2025supergpqascalingllmevaluation], and HLE [phan2025humanity0s] push the frontier further by demanding expert-level knowledge comprehension and deep reasoning.
Despite this progress, these benchmarks predominantly collect questions that already exist in textbooks or on websites. This paradigm of curating benchmarks at the question level suffers from three key limitations. First, questions or their variants may already appear in model training corpora. Even if a benchmark is initially uncontaminated, it risks exposure in future training, potentially distorting evaluation results. Second, each question typically targets a single knowledge concept, making it difficult to assess comprehensive understanding across multiple knowledge statements. Third, curating high-quality questions requires domain experts to design, annotate, and review them, incurring substantial cost and time.
@@
Figure 1: Overview of Encyclo-K: from statements to dynamically assembled evaluation questions.
Concretely, as illustrated in Figure 1, we extract standalone knowledge statements from authoritative textbooks and compose them into evaluation questions through random sampling at test time. This simple pipeline directly addresses all three limitations. For contamination resistance, we randomly sample and combine statements into questions at each test time. Even if individual statements appear in training data, their compositions form a combinatorial space too vast to memorize. For comprehensive assessment, each question aggregates 8–10 statements and asks models to identify which are correct. This requires joint understanding across diverse statements, going beyond what single-concept questions can probe. For cost reduction, correct statements are extracted directly from textbooks while incorrect ones are generated by reasoning models. Annotators only verify that each statement is self-contained and correctly formatted, requiring no expertise.
Extensive experiments on over 50 LLMs validate the effectiveness of Encyclo-K. Regarding dynamic evaluation, model rankings remain stable across five question sets generated with different random seeds, confirming that our combinatorial design resists memorization and enables reliable periodic dataset refresh. Regarding multi-statement comprehension, compared to single-concept questions in traditional benchmarks, models consistently exhibit over 20 percentage points lower accuracy on our multi-statement comprehensive understanding tasks, demonstrating the substantial challenges introduced by joint comprehension of multiple knowledge statements.
@@
  • We rethink benchmark construction by treating individual knowledge statements, rather than questions, as the atomic unit of curation. This shift enables diverse sourcing from authoritative materials, eliminates the need for expert annotation, and substantially reduces construction costs.
  • We compose statements into questions via random sampling, yielding two key properties: (1) Dynamic—the combinatorial space is too vast to memorize, with experiments across multiple random seeds confirming stable model rankings; (2) Multi-statement—each question requires joint comprehension of multiple statements, posing greater challenges than single-concept questions.
@@ 2.2 LLMs Evaluation Benchmarks
To ensure a robust evaluation of model generalizability and alleviate question leakage, we design two distinct mechanisms. First, Encyclo-K supports supports dynamic evaluation: statements within questions can be substituted, making the question content itself dynamically variable and enabling continuous generation of novel test samples to effectively address data contamination. Second, each question integrates multiple knowledge statements, emphasizing comprehensive understanding while ensuring that individual questions better represent the overall knowledge structure of a discipline.
3 Encyclo-K Dataset
@@ 5.1 Main Results
We first compare reasoning models against their chat counterparts within the same model family. As shown in Table 3, Qwen3-30B-A3B-Thinking (40.57%) outperforms Qwen3-30B-A3B-Instruct (36.72%) by 4 percentage points, and Qwen3-4B-Thinking (33.45%) exceeds Qwen3-4B-Instruct-2507 (28.60%) by approximately 5 percentage points. While this comparison suggests reasoning benefits, it conflates reasoning capability with potential differences in training data or optimization objectives between model variants.